# From Loops to Oops: Fallback Behaviors of Language Models Under Uncertainty

### v1

@@ B.2 Biographies Generations
Appendix C How Models Change Topics
@@
Inspired by this line of work, we analyzed how models change topics, and in particular, if memorized sequences emerge in these cases. We manually inspect such generations from multiple models, for both our TriviaFacts and BioGeneration datasets, using both greedy and temperature sampling. Using Infini-gram (Liu et al., 2024), we iterate over sliding windows of lengths 8 and 16 tokens to look for possible occurrences of verbatim recollection, which we manually inspect. We found no cases where a sequence longer than a few tokens repeats verbatim, and even then it comprises of common patterns, such as ‘‘public static void main’’.
@@
Appendix D Additional Results
@@
Figures 17 and 18 depict the scaling behavior of larger models on the TriviaFacts and Qampari datasets respectively. Figures 19 and 20 shows fallback trends during pretraining for Pythia-6.9B and 12B models in the former and OLMo-1B and 7B models in the latter. Figures 21 and 22 gives similar results when adding instruction finetuning to models on the TriviaFacts and Qampari datasets. In Figure 25 we ablate the prompt for instruction finetuned models to nudge them into abstaining rather than producing hallucinations, and Figure 26 uses in-context examples to do the same for pretrained models. Finally, Figures 27, 28, 29, 30 and 31 bring a breakdown of the results on the BioGeneration datasets with different levels of popularity.
D.1 Exclusion of Llama 2 13b Checkpoint
@@
D.2 Order of Facts in a Single Generation
@@
Additionally, in Section 6.1 we discuss the ordering of fact-labels within a single generation, as presented in Figure 7 for Pythia-12B model on TriviaFacts. Figures 32, 33, 34, 35, 36, 37, 38, 39, 40 and 41 give similar plots for other models.
@@
Figure 25: Instruction tuned models continue to hallucinate, even when told to prefer abstaining over non-factual responses. When given the prefix ‘‘Complete the following list with facts you are sure of, and stop when you cannot recall additional facts’’ to encourage abstaining instead of hallucinating, all instruction tuned models fail to utilize internal uncertainty estimations if exists and continue to hallucinate. The models are only slightly more inclined to abstain (by changing topics or producing EOS tokens). The results are given for the TriviaFacts dataset, with “(IDK)” marking the model with the modified prompt. The horizontal green line indicates the maximum number of correct answers possible.
@@
Figure 26: Larger models hallucinate more, even when shown how to abstain. When given the prefix from Figure 16 to encourage abstaining instead of hallucinating, larger Pythia and OLMo models continue to exhibit the same fallback behavior trends, with repetitions decreasing and hallucinations increasing as model size grows. The models are only slightly more inclined to abstain (by changing topics). Notably, while Llama 2 and Llama 3 models exhibit the predicted trends in the proportion of hallucinations, they are considerably more capable of abstaining when given the opportunity. Interestingly, Pythia-410M is able to change topics remarkably well, though when manually inspecting its outputs, we find that in almost all cases it produces a list of five facts, often with repetitions, and then continues to repeat this list indefinitely. The results are given for the TriviaFacts dataset, and the horizontal green line indicates the maximum number of correct answers possible.

### v2

@@ C.2 Biographies Generations
Appendix D How Models Change Topics
@@
Inspired by this line of work, we analyzed how models change topics, and in particular, if memorized sequences emerge in these cases. We manually inspect such generations from multiple models, for both our TriviaFacts and BioGeneration datasets, using both greedy and temperature sampling. Using Infini-gram (Liu et al., 2024a), we iterate over sliding windows of lengths 8 and 16 tokens to look for possible occurrences of verbatim recollection, which we manually inspect. We found no cases where a sequence longer than a few tokens repeats verbatim, and even then it comprises of common patterns, such as ‘‘public static void main’’.
@@
Appendix E Additional Results
@@
Figures 17 and 18 depict the scaling behavior of larger models on the TriviaFacts and Qampari datasets respectively. Figures 19 and 20 shows fallback trends during pretraining for Pythia-6.9B and 12B models in the former and OLMo-1B and 7B models in the latter. Figures 21 and 22 gives similar results when adding instruction finetuning to models on the TriviaFacts and Qampari datasets. In Figures 27 and 28 we ablate the prompt for instruction finetuned models to nudge them into abstaining rather than producing hallucinations on Triv and Qamp respectively, and Figures 25 and 26 uses in-context examples to do the same for base models. Section 5.1 shows how additional randomness in sampling reduces repetitions, and trades it with hallucinations (Figure 6. Figures 29, 30, 31, 32 and 33 provide additional evidence for these phenomena, repeating the experiments both with additional models, as well as using top-p sampling. Finally, Figures 35, 35, 37, 37 and 38 bring a breakdown of the results on the BioGeneration datasets with different levels of popularity.
E.1 Exclusion of Llama 2 13b Checkpoint
@@
E.2 Order of Facts in a Single Generation
@@
Additionally, in Section 6.1 we discuss the ordering of fact-labels within a single generation, as presented in Figure 8 for Pythia-12B model on TriviaFacts. Figures 40, 40, 42, 42, 44, 44, 46, 46, 48 and 48 give similar plots for other models.
@@
Figure 25: Larger models hallucinate more on TriviaFacts, even when shown how to abstain. When given the prefix from Figure 16 to encourage abstaining instead of hallucinating, larger Pythia and OLMo models continue to exhibit the same fallback behavior trends, with repetitions decreasing and hallucinations increasing as model size grows. The models are only slightly more inclined to abstain (by changing topics). Notably, while Llama 2 and Llama 3 models exhibit the predicted trends in the proportion of hallucinations, they are considerably more capable of abstaining when given the opportunity. Interestingly, Pythia-410M is able to change topics remarkably well, though when manually inspecting its outputs, we find that in almost all cases it produces a list of five facts, often with repetitions, and then continues to repeat this list indefinitely. The results are given for the TriviaFacts dataset, and the horizontal green line indicates the maximum number of correct answers possible.
Figure 26: Larger models hallucinate more on Qampari, even when shown how to abstain. When given the prefix from Figure 16 to encourage abstaining instead of hallucinating, larger Pythia and OLMo models continue to exhibit the same fallback behavior trends, with repetitions decreasing and hallucinations increasing as model size grows. The models are only slightly more inclined to abstain (by changing topics). Notably, while Llama 2 and Llama 3 models exhibit the predicted trends in the proportion of hallucinations, they are considerably more capable of abstaining when given the opportunity. Similarly to the phenomana observed in Figure 16, Pythia-410M is able to change topics more than expected, though upon manually inspection of its outputs, we find that in almost all cases it produces a list of just a few facts, often with repetitions, and then continues to repeat this list including the prompt indefinitely. The results are given for the Qampari dataset, and the horizontal green line indicates the maximum number of correct answers possible.
Figure 27: Instruction tuned models continue to hallucinate on TriviaFacts, even when instructed to generate only facts that the model is certain about. When given the prefix ‘‘Complete the following list with facts you are sure of, and stop when you cannot recall additional facts’’ to encourage abstaining instead of hallucinating, all instruction tuned models ( “(IDK)” marking the model with the modified prompt) fail to utilize internal uncertainty estimations if exists and continue to hallucinate. Some models are slightly more inclined to abstain (by changing topics or producing EOS tokens), while others attempt to compete the list even more than before resulting in additional hallucinations. The horizontal green line indicates the maximum number of correct answers possible.
Figure 28: Instruction tuned models continue to hallucinate on Qampari, even when instructed to generate only facts that the model is certain about. When given the prefix ‘‘Complete the following list with facts you are sure of, and stop when you cannot recall additional facts’’ to encourage abstaining instead of hallucinating, all instruction tuned models ( “(IDK)” marking the model with the modified prompt) fail to utilize internal uncertainty estimations if exists and continue to hallucinate. Some models are slightly more inclined to abstain (by changing topics or producing EOS tokens), while others attempt to compete the list even more than before resulting in additional hallucinations. The horizontal green line indicates the maximum number of correct answers possible.
Figure 29: Using larger $p$ with top-p sampling shifts fallback behavior from repetitions to hallucinations. An average of 5 completions from Pythia-12B on TriviaFacts for varying temperatures.
Figure 30: Using larger $p$ with top-p sampling shifts fallback behavior from repetitions to hallucinations. An average of 5 completions from Pythia-6.9B on TriviaFacts for varying temperatures.
Figure 31: Using larger $p$ with top-p sampling shifts fallback behavior from repetitions to hallucinations. An average of 5 completions from Pythia-2.8B on TriviaFacts for varying temperatures.
Figure 32: Increasing the sampling temperature shifts fallback behavior. An average of 5 completions from Pythia-6.9B on TriviaFacts for varying temperatures.
Figure 33: Increasing the sampling temperature shifts fallback behavior. An average of 5 completions from Pythia-2.8B on TriviaFacts for varying temperatures.